MARC:

By expressing cos2θ and sin2θ in terms of tanθ,solve the following eqn in the range 0≤θ≤360 . a) cos2θ−2sin2θ=2

MARC:

@mathmale

mathmale:

Sorry I didn't get to your question tonight (Thursday). Possibly tomorrow.

MARC:

sure,no prob.

celticcat:

use the identities cos 2θ = (1 - tan^2 θ ) / ( 1 + tan^2 θ ) and sin 2θ = 2 tan θ / ( 1 + tan^2 θ)

MARC:

Hmm,can u show me how u get that? I thought \(cos2\theta=cos^2\theta-sin^2\theta\) \(=2cos^2\theta-1\) \(=1-2sin^2\theta\) \(sin2\theta=2sin\theta~cos\theta\)

celticcat:

(1 - tan^2 θ ) / ( 1 + tan^2 θ ) - 2* 2 tan θ / ( 1 + tan^2 θ) = 2 multiply through by ( 1 + tan^2 θ):- 1 - tan^2 θ - 4 tan θ = 2(1 + tan^2 θ) 3 tan^2 θ + 4 tan θ + 1 = 0 solve this for tan θ - then you can find the required values of θ.

sillybilly123:

with \(t = \tan \theta\), you can grind these out: \(\sin 2 \theta = \dfrac{2t}{t^2 + 1} \qquad \triangle\) \(\cos 2 \theta = \dfrac{1-t^2}{t^2 + 1}\) Eg: \(\triangle\) follows from: \(\sin 2 \theta = 2 \sin \theta \cos \theta\) \( = 2 \tan \theta \cos^2 \theta\) \( = 2 \dfrac{\tan \theta }{\sec^2 \theta}\) \( = 2 \dfrac{\tan \theta }{\tan^2 \theta + 1} = \dfrac{2t}{t^2 + 1} \) Then have a play

MARC:

I think I know how to do...

MARC:

Thank you @celticcat @sillybilly123